# MSM_pseudo_CS

### v1

@@ 4 Product limit estimators
We propose a novel estimator for estimating the conditional state occupation probability $\Psi_{k|j}$. This estimator is defined in terms of marginal state occupation probabilities. Due to the tree structure of the multistate system, any state will be reached from the initial state 0 by a unique path. Define $\mathcal{S}^{k}$ as the set of states $\ell$ such that state $k$ is on the path from 0 to $\ell$. Then, the conditional quantity, $\Psi_{k|j}$, can be expressed as the ratio of two marginal functions given by
@@
where the numerator is the marginal occupation probability of occupying state $k$ or any of the other states that come after $k$ in the progressive system at time $t$, while the denominator is the marginal occupation probability of occupying state $j$ or any of the other states that come after $j$ in the system during an individual’s lifetime. We denote the resulting estimator defined by the ratio of marginal estimates of state occupation probabilities by $\widehat{\Psi}^{[2]}_{k|j}$. Since Equation (9) is a function of marginal state occupation probabilities, we begin with estimating these marginal probabilities by calculating the counting process and at-risk set using the method of Datta and Sundaram (2006), described in Section 2.1. Afterward, we obtain the Aalen-Johansen estimates of the marginal state occupation probabilities (see details below). Finally, $\widehat{\Psi}^{[2]}_{k|j}$ is obtained by the plug-in method.
where $\widehat{\pi}_{j^{\prime}}(0)=\{\widehat{Y}_{j^{\prime}}(0+)\}/\{\sum^{Q}_{j=1}\widehat{Y}_{j}(0+)\}$, and $\widehat{P}_{j^{\prime}j}(0,t)$ is the $j^{\prime}j$th element of $\widehat{\boldsymbol{P}}(0,t)$. Datta and Satten (2001) demonstrated that Equation (12) provides a consistent estimator for the state occupation probability, even in non-Markov models. This finding allows us to estimate the conditional state occupation probability without the Markov assumption using a product-limit approach similar to the Aalen-Johansen calculation.
Once we construct $\widehat{\Psi}^{[2]}_{k|j}$ using the product-limit approach, directly estimating the conditional entry distribution $F_{k|j}(t)$ becomes straightforward by substituting $\widehat{\Psi}^{[2]}_{k|j}$ into Equation (8).
@@ 4.1 A toy example
Figure 3: The estimates and 95% pointwise bootstrap confidence intervals for $\Psi_{5|1}$ based on simulated data from the seven-state illness death model using the two proposed methods. The dark solid line represents the true probability curve, the dark dotted line represents the estimated probability and the grey dashed lines denote the confidence intervals.
@@ 5 Bootstrap-based Confidence Intervals
The utilization of nonparametric regression in conjunction with the PAV step for formulating the estimators proposed in Section 3 and 4 imposes challenges for the asymptotic analysis for these estimators. To quantify uncertainty, a bootstrap-based procedure may provide a practical alternative. However, Faraway and Jhun (1990) pointed out that a direct application of the bootstrap will generally lead to invalid results for the “smoothed” nonparametric estimators since it cannot capture the bias component of the smoothed estimator.
@@
For $10<\alpha<$ 1, let $\widehat{\Delta}_{\alpha}(t)$ be the $(1-\alpha)$ bootstrap percentile of the distribution of
$\big{|}\sin^{-1}\big{(}\sqrt{\widehat{\theta}^{*}_{k|j}(t;h)}\big{)}-\sin^{-1}\big{(}\sqrt{\widehat{\theta}_{k|j}(t;\tilde{h})}\big{)}\big{|}$
where $\widehat{\theta}^{*}_{k|j}(t;h)$ is calculated using the bandwidth $h$ from the original estimation but is based on the bootstrap sample, while $\widehat{\theta}_{k|j}(t;\tilde{h})$ uses the new bandwidth $\tilde{h}$ but is based on the original sample. The $(1-\alpha)\times 100\%$ pointwise CI for the temporal function $\theta_{k|j}(t)$ is given by
[…]
$\quad\sin^{2}\Big{\{}\min\Big{(}\frac{\pi}{2},\sin^{-1}\Big{(}\sqrt{\widehat{\theta}_{k|j}(t;h)}\Big{)}+\widehat{\Delta}_{1-\alpha/2}\Big{)}\Big{\}}\Big{]}.$
The construction of the confidence interval described above is similar to Datta and Sundaram (2006). In the subsequent section, we conduct a Monte Carlo study to examine the behavior of the smoothed bootstrap confidence intervals.
6 Simulation study
We performed extensive simulation studies to evaluate the performance of the proposed methods based on: (i) fractional risk sets (FRE), and (ii) the ratio of product limit estimators (PLE) in calculating the conditional estimators. We considered two separate multistate systems - the five-state illness-death model shown in Figure 1, and the seven-state multistate model shown in Figure 4. For each multistate system, the exact transition times were generated from a common lognormal hazard. We considered sample sizes of 100, 200, 500, and 1000.
@@ 6.1 Five-State Illness-Death Model
Earlier, in Section 3.1, we introduced the five-state illness-death model. For our simulation experiment, our primary interest lies in estimating the probability of occupying state 3 having visited state 1, denoted by $\Psi_{3|1}$. We also study the performance of the estimators for the subdistribution function $\Psi_{3|1}(t)$, and the future entry distribution for state 3 given the prior occupation of state 1, denoted by $F_{3|1}(t)$. Our data simulation process initiates with each individual starting in state 0 at time 0, with potential transitions to either state 1 or state 2. Specifically, we simulate the scenario where 60% of individuals follow the $0\rightarrow 1$ transition path, while the remaining 40% follow the $0\rightarrow 2$ transition path. The transition out of state 0 is governed by an independent Bernoulli distribution. For those individuals transitioning to state 1, we further simulate their progression, with 60% transitioning to state 3 and 40% to state 4. This transition from state 1 is also controlled by a separate independent Bernoulli distribution. Following this simulation scheme, the true value of $\Psi_{3|1}$ is 0.6.
@@ 6.2 Seven-State Multistate Model
Here, we consider the more complex multistate system shown in Figure 4. Primarily, we are interested in estimating the probability of occupying State 5 having visited State 1, denoted by $\Psi_{5|1}$. From Eqn 6, we have $\Psi_{5|1}=\Psi_{3|1}\times\Psi_{5|3}$. In addition to studying the performance of the estimators for $\Psi_{5|1}$, we will also study the performance of the estimators for $\Psi_{5|3}$, along with the subdistribution functions; $\Psi_{5|1}(t)$ and $\Psi_{5|3}(t)$, and the future entry distributions; $F_{5|1}(t)$, and $F_{5|3}(t)$. For the data simulation, all individuals start in state 0 at time 0, and they have two possible paths for transition: either to state 1 or state 2. We generate the data as follows: 80% of the individuals follow the $0\rightarrow 1$ path, while the remaining 20% follow the $0\rightarrow 2$ path. Among those who transitioned to state 1, there are two possible paths: 70% transition to state 3, and the remaining 30% transition to state 4. Moreover, among those that visited state 3, 60% proceeded to state 5, while the remaining 40% transitioned to state 6. Note that, the respective transitions out of state 0, state 1, and state 3 are controlled by three separate Bernoulli distributions. Following this simulation approach, the true value of $\Psi_{5|1}$ is 0.6.
@@ 6.3 Simulation results
We evaluated the performance of the estimator of the conditional state occupation probability $\Psi_{k|j}$ using absolute bias. For the global assessment of the estimators of the subdistribution functions, $\Psi_{k|j}(t)$, and the future state entry time distribution, $F_{k|j}(t)$, we employed the mean absolute distance (MAD), defined as follows:
[…]
@@
Table 2 shows the simulation results for the seven-state multistate system where the true transition times are generated from the lognormal hazard. For each quantity of interest estimated by either the FRE or PLE approach, the performance of the estimators appears to improve with increasing sample size. Under this system, we observe that the estimators based on the FRE method slightly outperform the estimators based on the PLE method. As pointed out in Lan and Datta (2010), a possible explanation of the superiority of the FRE method over the PLE method especially for later states in the system is that the PLE-based estimation of the functions for the later states depends on the estimation done in prior states, therefore, the estimation error may propagate along the path.
Table 1: Absolute bias ($\Psi_{3|1}$) and MAD ($\Psi_{3|1}(t)$ $F_{3|1}(t)$) for the scenario where the true transition times for the five-state illness-death model were generated from the lognormal distribution. Inspection time for each individual was generated from (i) uniform and (ii) Weibull distribution. The results are shown by the methods proposed: (a) FRE, and (b) PLE. The results are based on a Monte Carlo sample size of 1000.
$\Psi_{3|1}$ $\Psi_{3|1}(t)$ $F_{3|1}(t)$
Inspection time $N$ FRE PLE FRE PLE FRE PLE
100 0.057 0.065 0.055 0.059 0.065 0.061
200 0.039 0.041 0.040 0.042 0.049 0.046
500 0.026 0.025 0.029 0.029 0.035 0.032
Uniform 1000 0.019 0.018 0.023 0.022 0.027 0.025
100 0.047 0.060 0.053 0.065 0.082 0.087
200 0.036 0.045 0.041 0.047 0.070 0.072
500 0.028 0.030 0.029 0.032 0.054 0.054
Weibull 1000 0.025 0.025 0.023 0.024 0.046 0.045
Table 2: Absolute bias ($\Psi_{5|1}\ \&\ \Psi_{5|3}$) and MAD ($\Psi_{5|1}(t),\ \Psi_{5|3}(t),\ F_{5|1}(t)\ \&\ F_{5|3}(t)$) for the scenario where the true transition times for the seven-state illness-death model were generated from the lognormal distribution. Inspection time for each individual was generated from (i) uniform and (ii) Weibull distribution. The results are shown by the methods proposed: (a) FRE, and (b) PLE. The results are based on a Monte Carlo sample size of 1000.
$\Psi_{5|1}$ $\Psi_{5|3}$ $\Psi_{5|1}(t)$ $\Psi_{5|3}(t)$ $F_{5|1}(t)$ $F_{5|3}(t)$
Inspection time $N$ FRE PLE FRE PLE FRE PLE FRE PLE FRE PLE FRE PLE
100 0.052 0.060 0.059 0.068 0.044 0.049 0.049 0.056 0.060 0.059 0.057 0.059
200 0.037 0.043 0.042 0.050 0.035 0.039 0.039 0.045 0.049 0.048 0.045 0.048
500 0.024 0.027 0.027 0.031 0.026 0.029 0.028 0.034 0.038 0.040 0.033 0.040
Uniform 1000 0.018 0.019 0.021 0.022 0.021 0.025 0.022 0.029 0.030 0.035 0.025 0.035
100 0.054 0.056 0.068 0.073 0.045 0.046 0.054 0.058 0.074 0.072 0.072 0.072
200 0.036 0.044 0.040 0.048 0.034 0.034 0.038 0.040 0.060 0.060 0.058 0.060
500 0.027 0.031 0.030 0.034 0.026 0.023 0.028 0.028 0.049 0.048 0.047 0.048
Weibull 1000 0.022 0.023 0.025 0.027 0.023 0.018 0.022 0.022 0.044 0.041 0.038 0.041
@@ 6.3.1 Coverage of smoothed bootstrap confidence intervals
For each multistate system considered in this simulation study, we evaluate the performance of the smoothed bootstrap confidence intervals. To obtain the confidence intervals, we used 1000 bootstrap resamples from the original data to calculate the bootstrap percentiles. The reported empirical coverage probabilities are based on 1000 Monte Carlo iterations at a nominal level of 0.05. Table 3 shows the empirical coverage probabilities for the estimators of $\Psi_{3|1}(t)$ based on the five-state illness death model. The coverage probabilities for the estimators of $\Psi_{5|1}(t)$ based on the seven-state illness death model are shown in Table 4. The confidence intervals for the temporal functions of interest are calculated at five fixed time points. The results are shown for the scenario where the random inspection time for each individual is generated from a uniform distribution. From Tables 3 and 4, the coverage probabilities for the estimator based on both FRE and PLE approach are reasonable.
Table 3: Coverage probabilities for the estimator of $\Psi_{3|1}(t)$ for the two proposed methods; (i) FRE, and (ii) PLE. The confidence intervals are based on 1000 bootstrap samples at a nominal level of 0.05, and the coverage results are based on 1000 Monte Carlo iterations.
FRE PLE
Sample size Sample size
Time 100 200 500 1000 100 200 500 1000
0.75 0.985 0.953 0.928 0.905 0.940 0.971 0.987 0.993
1.25 0.945 0.956 0.977 0.976 0.906 0.937 0.968 0.984
1.75 0.975 0.970 0.977 0.976 0.981 0.982 0.992 0.995
2.25 0.989 0.989 0.978 0.989 0.995 0.996 0.997 0.999
2.75 0.996 0.995 0.997 0.985 0.999 0.999 0.997 0.999
Table 4: Coverage probabilities for the estimator of $\Psi_{5|1}(t)$ for the two proposed methods; (i) FRE, and (ii) PLE. The confidence intervals are based on 1000 bootstrap samples at a nominal level of 0.05, and the coverage results are based on 1000 Monte Carlo iterations.
FRE PLE
Sample size Sample size
Time 100 200 500 1000 100 200 500 1000
2.00 0.858 0.873 0.840 0.813 0.929 0.955 0.979 0.985
3.00 0.853 0.851 0.797 0.764 0.976 0.983 0.988 0.968
4.00 0.970 0.957 0.945 0.935 0.994 0.992 0.996 0.978
5.00 0.994 0.994 0.986 0.984 0.998 0.999 0.997 0.994
6.00 0.999 1.000 0.997 0.998 0.999 0.999 0.999 1.000
7 Breast cancer example (EORTC-trial 10854)
To further illustrate the two proposed estimation methods, we re-analyzed data on 2793 breast cancer patients who participated in a trial conducted by the European Organization for Research and Treatment of Cancer (EORTC-trial 10854). The trial which involved patients with early breast cancer, aimed to compare whether surgery followed by a short intensive course of perioperative chemotherapy yields better treatment than surgery alone. A comprehensive description of the study is provided in Clahsen et al. (1996), and further findings are reported in Van der Hage et al. (2001).
Figure 4 describes the multistate system employed to illustrate our proposed methods. This system comprises nine states, including event-free and alive after surgery (state 0), loco-regional relapse following surgery (state 1), distant metastasis following surgery (state 2), simultaneous occurrence of loco-regional relapse and distinct metastasis following surgery (state 3), death without the occurrence of loco-regional relapse or distinct metastasis following surgery (state 4), distant metastasis following occurrence of loco-regional relapse after surgery (state 5), death following occurrence of loco-regional relapse after surgery (state 6), loco-regional relapse following occurrence of distant metastasis after surgery (state 7), and death following occurrence of distant metastasis after surgery (state 8). The time origin for the analysis was set as the time of surgery, meaning all individuals commenced at state 0. Transition times in the original dataset were right-censored, with follow-up times ranging from 12 to 6,756 days, and a median of 3,817 days.
To obtain current status data, the inspection time for each patient was randomly generated from a uniform distribution with a lower bound of 0 and an upper bound of 6,756 days, which is the maximum follow-up time observed across all patients in the original data. The inspection times were independent of the progression between the event states. We defined a patient’s inspection state as the state occupied at their inspection time. If the patient was right-censored, the inspection state was the most recent state they occupied before censoring. To highlight the structure of the current status data from the multistate system, we provide information for five randomly selected individuals in the supplementary material. This cross-sectional version of the data poses a severe type of censoring since we would only know the state occupied by each patient at a random inspection time rather than their transition history up to a censoring time.

### v2

@@ 4 Product limit estimators
We introduce a novel estimator for the conditional state occupation probability $\Psi_{k|j}$. This estimator is expressed in terms of marginal state occupation probabilities, which leverage the tree structure of the multistate system. In such a system, any state is reached from the initial state, 0, through a unique path. Define $\mathcal{S}^{k}$ as the set of states $\ell$ such that state $k$ is on the path from 0 to $\ell$. Then, the conditional quantity, $\Psi_{k|j}$, can be expressed as the ratio of two marginal functions given by
@@
where the numerator is the marginal occupation probability of occupying state $k$ or any of the other states that come after $k$ in the progressive system at time $t$, while the denominator is the marginal occupation probability of occupying state $j$ or any of the other states that come after $j$ in the system during an individual’s lifetime. The resulting estimator, denoted by $\widehat{\Psi}^{[2]}_{k|j}$, is constructed from the ratio of these marginal state occupation probabilities.
To construct $\widehat{\Psi}^{[2]}_{k|j}$, we begin by estimating the marginal state occupation probabilities using the product-limit (Aalen-Johansen) formula. This method involves first estimating the counting process and at-risk set, followed by the calculation of the marginal state occupation probabilities. For completeness, we provide the development of these nonparametric estimators under current status censoring in Appendix A, following the methodology of Datta and Sundaram [10]. Once these marginal estimates are obtained, the estimator $\widehat{\Psi}^{[2]}_{k|j}$ is derived through a straightforward plug-in procedure.
Following the construction of $\widehat{\Psi}^{[2]}_{k|j}$ using the product-limit approach, the estimation of the conditional entry distribution $F_{k|j}(t)$ becomes straightforward by the substitution of $\widehat{\Psi}^{[2]}_{k|j}$ into (6).
@@ 4.1 A toy example
Figure 3: Estimated probability curves for $\Psi_{5|1}(t)$ using the two proposed methods, along with 95% pointwise bootstrap confidence intervals, based on simulated data from a seven-state model. For $N=1000$ individuals, true transition times were generated from a lognormal distribution, while inspection times followed a uniform distribution.
@@ 5 Bootstrap-based Confidence Intervals
The utilization of nonparametric regression in conjunction with the PAV step for formulating the estimators proposed in Section 3 and 4 imposes challenges for the asymptotic analysis for these estimators. To quantify uncertainty, a bootstrap-based procedure may provide a practical alternative. However, Faraway and Jhun [12] pointed out that a direct application of the bootstrap will generally lead to invalid results for the “smoothed” nonparametric estimators since it cannot capture the bias component of the smoothed estimator.
Following guidelines from existing work on the bootstrap for nonparametric regression estimators [19, 27], we describe a smoothed bootstrap procedure for estimating the pointwise confidence intervals for our estimators. Let $\theta(t)$ denote either $\Psi_{k\mid j}(t)$ or $F_{k\mid j}(t)$. First, using random sampling with replacement, we generate bootstrap censored times $\{C^{*}_{i};\ 1,...,n\}$ from the original data $\{C_{i};\ 1,...,n\}$. For the $i$th observation, we then generate the corresponding state $S^{*}_{i}(C^{*}_{i})$ from the distribution given by
In other words, we generate $S^{*}_{i}(C^{*}_{i})$ from a smoothed distribution of the states whose inspection times were close to $C^{*}_{i}$. Smoothing reduces variance by borrowing strength across nearby estimates, but at the potential cost of increased bias. The bias is mitigated by selecting an appropriate bandwidth using data-driven approaches (see Section A.1), as excessively large bandwidths can oversmooth, distorting the probability estimates. Following the theory of the smoothed bootstrap, the bandwidth $\tilde{h}=h^{\omega}$ needs to be larger than the bandwidth $h$ selected in the estimation based on the original data. We recommend setting $\omega=0.8$ if $0<h<1$, and $\omega=1.2$ if $h>1$.
Because $\theta(t)$ is a probability-valued function and may lie near 0 or 1, we apply the variance-stabilizing transformation, $g(x)=\sin^{-1}(\sqrt{x})$, which produces a more symmetric and stable sampling distribution. The bootstrap distribution of
@@
$\Big[g^{-1}\Big\{$ $\max\Big(0,g\Big(\widehat{\theta}(t;h)\Big)-\widehat{\Delta}_{1-\alpha/2}(t)\Big)\Big\},\ g^{-1}\Big\{\min\Big(\frac{\pi}{2},g\Big(\widehat{\theta}(t;h)\Big)+\widehat{\Delta}_{1-\alpha/2}(t)\Big)\Big\}\Big],$
where $g^{-1}(x)=\sin^{2}(x)$. The truncation at 0 and $\pi/2$ ensures that the inverse transformation yields valid probability bounds in $[0,1]$. The construction of the confidence interval described above is similar to Datta and Sundaram [10]. In Section 7, we conduct a Monte Carlo study to examine the behavior of the smoothed bootstrap confidence intervals.
6 Testing of covariate effects via pseudo-value regression
We use pseudo-value regression to test the effect of baseline covariates, $\mathbf{Z}$, on the distribution of state entry times conditional on a prior state visit. Originally proposed by Andersen et al. [2] for multistate models with right censoring, this approach has recently been validated by Anyaso-Samuel et al. [4] for multistate current status data.
Let $\widehat{\theta}(t)$ denote the marginal estimator of interest, which, in our context, may represent either $\Psi_{k|j}(t)$ or $F_{k|j}(t)$. The jackknife pseudo-values are defined as:
$Y_{i}(t)=n\widehat{\theta}(t)-(n-1)\widehat{\theta}_{-i}(t),$ (8)
where $\widehat{\theta}_{-i}(t)$ is calculated by omitting the $i$th subject from the estimation. Each pseudo-value $Y_{i}(t)$ quantifies the individual contribution of the $i$th subject to the marginal estimator and can be modeled approximately linearly in regression analyses.
To evaluate covariate effects, pseudo-values serve as the response variables in regression models. Specifically, we compute pseudo-values at ten equally spaced points across the observed event-time range. Generalized estimating equations (GEEs) with an independence working correlation structure are then used for inference regarding covariate effects.
We provide detailed steps for implementing pseudo-value regression in Appendix B. Extensive simulation studies, presented in Section A of the Web Supplementary material, validate the performance of this procedure.
7 Simulation studies
We performed extensive simulation studies to evaluate the performance of the proposed methods based on: (i) fractional risk sets (FRE), and (ii) the ratio of product limit estimators (PLE) for estimating the conditional quantities of interest. This section focuses on settings without covariate adjustment; results for scenarios with covariates are presented in Section A of the Web Supplementary Material. We considered two separate multistate systems: the five-state illness-death model shown in Figure 1, and the seven-state multistate model shown in Figure 2. For each multistate system, the exact transition times were generated from a common lognormal hazard. We considered sample sizes of 100, 200, 500, and 1000.
@@ 7.1 Five-State Illness-Death Model
Earlier, in Section 3.1, we introduced the five-state illness-death model. For our simulation experiment, our primary interest lies in estimating the probability of occupying state 3 having visited state 1, denoted by $\Psi_{3|1}$. We also study the performance of the estimators for the subdistribution function $\Psi_{3|1}(t)$, and the entry distribution for state 3 given the prior occupation of state 1, denoted by $F_{3|1}(t)$. Our data simulation process initiates with each individual starting in state 0 at time 0, with potential transitions to either state 1 or state 2. Specifically, we simulate the scenario where 60% of individuals follow the $0\rightarrow 1$ transition path, while the remaining 40% follow the $0\rightarrow 2$ transition path. The transition out of state 0 is governed by an independent Bernoulli distribution. For those individuals transitioning to state 1, we further simulate their progression, with 60% transitioning to state 3 and 40% to state 4. This transition from state 1 is also controlled by a separate independent Bernoulli distribution. Following this simulation scheme, the true value of $\Psi_{3|1}$ is 0.6.
@@ 7.2 Seven-State Multistate Model
Here, we consider the more complex multistate system shown in Figure 2. Primarily, we are interested in estimating the probability of occupying State 5 having visited State 1, denoted by $\Psi_{5|1}$. From (4), we have $\Psi_{5|1}=\Psi_{3|1}\times\Psi_{5|3}$. In addition to studying the performance of the estimators for $\Psi_{5|1}$, we will also study the performance of the estimators for $\Psi_{5|3}$, along with the subdistribution functions; $\Psi_{5|1}(t)$ and $\Psi_{5|3}(t)$, and the condition state entry distributions; $F_{5|1}(t)$, and $F_{5|3}(t)$. For the data simulation, all individuals start in state 0 at time 0, and they have two possible paths for transition: either to state 1 or state 2. We generate the data as follows: 80% of the individuals follow the $0\rightarrow 1$ path, while the remaining 20% follow the $0\rightarrow 2$ path. Among those who transitioned to state 1, there are two possible paths: 70% transition to state 3, and the remaining 30% transition to state 4. Moreover, among those that visited state 3, 60% proceeded to state 5, while the remaining 40% transitioned to state 6. Note that, the respective transitions out of state 0, state 1, and state 3 are controlled by three separate Bernoulli distributions. Following this simulation approach, the true value of $\Psi_{5|1}$ is 0.6.
@@ 7.3 Simulation results
We evaluated the performance of the estimator of the conditional state occupation probability $\Psi_{k|j}$ using absolute bias. For the global assessment of the estimators of the subdistribution functions, $\Psi_{k|j}(t)$, and the conditional state entry time distribution, $F_{k|j}(t)$, we employed the mean absolute distance (MAD), defined as follows:
$\text{MAD}=\mathbb{E}\int|\widehat{\theta}^{\text{Emp}}(x)-\widehat{\theta}(x)|\ d\theta_{n}(x),$ (9)
@@
Table 2 shows the simulation results for the seven-state multistate system where the true transition times are generated from the lognormal hazard. For each quantity of interest estimated by either the FRE or PLE approach, the performance of the estimators appears to improve with increasing sample size. Under this system, we observe that the estimators based on the FRE method slightly outperform the estimators based on the PLE method. As pointed out in [18], a possible explanation of the superiority of the FRE method over the PLE method especially for later states in the system is that the PLE-based estimation of the functions for the later states depends on the estimation done in prior states, therefore, the estimation error may propagate along the path.
@@ 7.3.1 Coverage of smoothed bootstrap confidence intervals
For each multistate system considered in this simulation study, we evaluate the performance of the smoothed-bootstrap pointwise confidence intervals. Confidence intervals were obtained using 1000 bootstrap resamples, and empirical coverage probabilities were computed from 1000 Monte Carlo iterations at a nominal level of 0.05. Table 3 reports coverage for the estimators of $\Psi_{3\mid 1}(t)$ under the five-state illness–death model, while Table 4 reports coverage for $\Psi_{5\mid 1}(t)$ under the seven-state model. The results correspond to the setting where inspection times are generated from a uniform distribution.
Because current-status inference at time $t$ is driven primarily by subjects inspected near $t$, coverage can be unstable near boundary regions where the inspection-time density (and hence the effective sample size) is small. We therefore report coverage at five fixed time points chosen away from the boundaries of the inspection-time support. Overall, FRE yields coverage close to the nominal level with mild deviations, whereas PLE tends to be conservative, with coverage often exceeding 0.95. The conservativeness is more pronounced for $\Psi_{5\mid 1}(t)$, reflecting the reduced information available for a state deep in the progressive tree, particularly at later times.
8 Application to emulated current status data from a real-world study
To illustrate our proposed estimators, we used data from a large randomized clinical trial conducted by the European Organization for Research and Treatment of Cancer (EORTC trial 10854), involving 2,793 patients with early-stage breast cancer. The trial compared outcomes between patients receiving surgery alone versus those receiving surgery plus perioperative chemotherapy. Further details are available in Clahsen et al. [6] and Van der Hage et al. [13]. For this analysis, we constructed a current status version of the dataset to reflect scenarios where each patient is observed only once at a random inspection time. Such designs are relevant in large-scale population studies and clinical follow-up programs where logistical or ethical considerations limit repeated assessments. Similar approaches have been used in recent work [33] to evaluate methodological robustness under severe censoring. Our goal is to examine how well the proposed estimators perform under these realistic but challenging conditions.
Figure 4 outlines the progressive nine-state system used in this analysis, with surgery as the time origin (state 0). Subsequent states represent key clinical events, including loco-regional recurrence (state 1), distant metastasis (state 2), simultaneous relapse and metastasis (state 3), and various death states depending on the patient’s prior path. Follow-up times ranged from 12 days to 18.5 years, with a median of 10.5 years.
To emulate current status data, each patient was assigned a single inspection time drawn uniformly between 0 and 18.5 years. The state occupied at that inspection time was recorded. If the patient was censored before the assigned inspection time, the most recently observed state was used. This framework represents a severe form of censoring, where only the state at a single, possibly late, time point is known. Table 5 summarizes the distribution of observed states at inspection. As expected, a large proportion of patients remained in the initial state (alive post-surgery) at inspection.